OpenStudy (anonymous):

Evaluate the given limit by first recognizing the indicatd sum as a Riemann sum associated with a regular partiion of [0,1] and then evaluating the corresponding integral. (a) lim (1+2+3+......+n)/n^2 as n approaches to positive infinity.

OpenStudy (anonymous):

we can evaluate this limit easily without doing that \[\sum_{k=1}^nk=\frac{n(n+1)}{2}\] so \[\frac{1}{n^2}\sum_{k=1}^nk=[frac{1}{n^2}\frac{n(n+1)}{2}=\frac{n^2+n}{2n^2}=\frac{1}{2}+\frac{1}{n}\] take limit as n goes to infinity and get \[\frac{1}{2}\]

OpenStudy (anonymous):

stupid typo \[\frac{1}{n^2}\sum_{k=1}^nk=\frac{1}{n^2}\frac{n(n+1)}{2}=\frac{n^2+n}{2n^2}=\frac{1}{2}+\frac{1}{n}\]

OpenStudy (anonymous):

or you could recognize this as \[\int_0^1 xdx=\frac{1}{2}\]

OpenStudy (anonymous):

how can you tell it is x?